OpenStudy (anonymous):

Perform the indicated operation. will some2 help me please :(

OpenStudy (anonymous):

whats the question

OpenStudy (anonymous):

what is the operation?

OpenStudy (anonymous):

\[\frac{ y }{ y-5 }+\frac{ 5 }{ 5-y}\]

OpenStudy (anonymous):

ok you can make the denominators the same by rewriting as \[\frac{y}{y-5}-\frac{5}{y-5}\]

OpenStudy (anonymous):

then subtract in the numerator and get \[\frac{y-5}{y-5}\] which is equal to \(1\)

OpenStudy (anonymous):

can u help me again